OpenStudy (anonymous):

\int\limits_{0}^{1} \int\limits_{0}^{X} (x^2 + 4xy) dy dx

geerky42 (geerky42):

Like this? \[\Huge \displaystyle \int\limits\limits_{0}^{1} \int\limits\limits_{0}^{x} (x^2 + 4xy) dy dx\]

OpenStudy (anonymous):

oh gosh this is a mess

OpenStudy (anonymous):

yes just like that